A particle is thrown vertically upwards from the ground. If the ratio of time of ascent to time of descent is √2/√3, then retardation produced due to resistive force acting on the particle is
 (a) 4 m/s^2
 (b) 2 m/s^2
 (c) 3 m/s^2
 (d) 6 m/s^2

Grade:11

1 Answers

Khimraj
5 years ago
ta/td = \sqrt{(g-a)/(g+a)}
(ta/td )2 = (g-a)/(g+a)
2/3 = (10-a)/(10+a)
20 + 2a = 30 – 3a
5a = 10
so a = 2 m/s2
Hope it clears........................................................................